Forrest M. Hoffman

Computational Earth Sciences Group Oak Ridge National Laboratory (ORNL) Oak Ridge, Tennessee 37831-6016 865-576-7680

                               forrest@climatemodeling.org

Education

University of Tennessee, Knoxville Physics B.S. 1991
University of Tennessee, Knoxville Physics M.S. 2004

Employment

Computational Climate Scientist, joint appointments in the Computer Science & Mathematics Division and the Environmental Sciences Division, Oak Ridge National Laboratory, Oak Ridge, Tennessee (November 1993–present).

Scientific Programmer/Analyst, Department of Geological Sciences, University of Tennessee, position at Environmental Sciences Division, Oak Ridge National Laboratory, Oak Ridge, Tennessee (July 1992– November 1993).

Knowledge Engineer, Automated Sciences Group, Inc., position at Environmental Sciences Division, Oak Ridge National Laboratory, Oak Ridge, Tennessee, (May 1989–June 1992).

Researcher Intern, Oak Ridge Associated Universities (ORAU), position at Environmental Sciences Division, Oak Ridge National Laboratory, Oak Ridge, Tennessee (September 1988–May 1989).

Observer/Telescope Operator, High Altitude Observatory, National Center for Atmospheric Research, position at Mauna Loa Solar Observatory, Hilo, Hawai’i (1987–September 1988).

Selected Publications

Silvia Kloster, Natalie M. Mahowald, James T. Randerson, Peter E. Thornton, Forrest M. Hoffman, Samuel Levis, Peter J. Lawrence, Johan J. Feddema, Keith W. Oleson, and David M. Lawrence. Fire dynamics during the 20th century simulated by the Community Land Model. Biogeosci., 7(6):1877–1902, June 2010. doi:10.5194/bg-7-1877-2010.

James T. Randerson, Forrest M. Hoffman, Peter E. Thornton, Natalie M. Mahowald, Keith Lindsay, Yen-Huei Lee, Cynthia D. Nevison, Scott C. Doney, Gordon Bonan, Reto Stöckli, Curtis Covey, Steven W. Running, and Inez Y. Fung. Systematic assessment of terrestrial biogeochemistry in coupled climate-carbon models. Global Change Biol., 15(10):2462–2484, October 2009. ISSN 1365-2486. doi:10.1111/j.1365-2486.2009.01912.x.

David J. Erickson III, Richard T. Mills, Jay Gregg, T. J. Blasing, Forrest M. Hoffman, Robert J. Andres, Matthew Devries, Z. Zhu, and S. R. Kawa. An estimate of monthly global emissions of anthropogenic CO_{2}: Impact on the seasonal cycle of atmospheric CO_{2}. J. Geophys. Res., 113(G1):G01023, March 2008. doi:10.1029/2007JG000435.

Forrest M. Hoffman, Curtis C. Covey, Inez Y. Fung, James T. Randerson, Peter E. Thornton, Yen-Huei Lee, Nan A. Rosenbloom, Reto C. Stöckli, Steven W. Running, David E. Bernholdt, and Dean N. Williams. Results from the Carbon-Land Model Intercomparison Project (C-LAMP) and availability of the data on the Earth System Grid (ESG). J. Phys.: Conf. Ser., 78(1):012026, December 2007. doi: 10.1088/1742-6596/78/1/012026.

Robert E. Dickinson, Keith W. Oleson, Gordon Bonan, Forrest Hoffman, Peter Thornton, Mariana Vertenstein, Zong-Liang Yang, and Xubin Zeng. The Community Land Model and its climate statistics as a component of the Community Climate System Model. J. Clim., 19(11):2302–2324, June 2006. doi: 10.1175/JCLI3742.1.

Forrest M. Hoffman, William W. Hargrove, David J. Erickson, and Robert J. Oglesby. Using clustered climate regimes to analyze and compare predictions from fully coupled general circulation models. Earth Interact., 9(10):1–27, August 2005. doi:10.1175/EI110.1.

Michael A. White, Forrest Hoffman, William W. Hargrove, and Ramakrishna R. Nemani. A global framework for monitoring phenological responses to climate change. Geophys. Res. Lett., 32(4):L04705, February 2005. doi:10.1029/2004GL021961.